OpenStudy (anonymous):

Jessica deposits $5,000 at the end of each year in an account earning 2.45% interest, compounded annually. What is the future value of this annuity after 5 years of investing? $25,246.20 $13,127.69 $52,510.76 $26,255.38

OpenStudy (anonymous):

I got $26,255.38 as my answer but I'm not sure if it's right

OpenStudy (anonymous):

I can check if you'd like.

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

KK

OpenStudy (anonymous):

5000 times 5 to start, because she adds this much at the end of each year. This gives you 25,000. Now you can eliminate B and C. Next, plug everything into the equation A=P*e^rt. This should give you A. $25, 246.20.

OpenStudy (anonymous):

A

OpenStudy (anonymous):

hope I have helped!!\[BYE\]

OpenStudy (tkhunny):

Time = 0 Years: 0 Time = 1 Years: 5000 Time = 2 Years: 5000(1.0245) + 5000 Time = 3 Years: 5000(1.0245)^2 + 5000(1.0245) + 5000 Time = 4 Years: 5000(1.0245)^3 + 5000(1.0245)^2 + 5000(1.0245) + 5000 Time = 4 Years: 5000(1.0245)^4 + 5000(1.0245)^3 + 5000(1.0245)^2 + 5000(1.0245) + 5000 Algebra: 5000(1 + 1.0245 + 1.0245^2 + 1.0245^3 + 1.0245^4) = \(5000\dfrac{1.0245^{5} - 1}{1.0245 - 1}\) = 5000(5.251076408) = $26,255.38 If you learn to build it, you need never be anything but confident.

OpenStudy (anonymous):

thanks a lot for the explanation it really helped!